# Diff-Shadow: Global-guided Diffusion Model for Shadow Removal

### v1

@@ 3.3 Global-guided Samping Strategy
$\small\frac{1}{\sum_{i}W^{d_{i}}}\sum_{i=1}^{N_{p}}\left(W^{d_{i}}\!\cdot\!\epsilon_{\theta}\left(\mathbf{x}_{t}^{(i)},{\mathbf{x}_{L}}^{(i)},{\mathbf{x}_{G}},t\right)\right),$ (8)
@@
3.4 Optimization
The Diff-Shadow involves two parallel UNet networks operating the local patch optimization and the global shadow-free image reconstruction. The objective function includes the following two items: the diffusive objective function $L_{\text{diff}}$ and the global loss function $L_{\text{global}}$.
$\small\mathcal{L}_{\text{total }}=\mathcal{L}_{\text{diff }}+\lambda\mathcal{L}_{\text{global}},$ (9)
where $\lambda$ is a hyper-parameter to balance the contributions of the two losses. We empirically set it to 1.
The Diffusive Objective Function. Refer to [17], the $\mathcal{L}_{\text{diff}}$ is defined as:
$\small\mathcal{L}_{\text{diff}}=\mathbb{E}_{\mathbf{x}_{0},t,\boldsymbol{\epsilon}}\left\|\boldsymbol{{\epsilon}}_{t}-\boldsymbol{\epsilon}\right\|_{F}^{2}.$ (10)
@@ 3.4 Optimization
$\small\mathcal{L}_{\text{global}}=\mathbb{E}_{\mathbf{{x}_{0}}^{\prime},\mathbf{{x}_{m}}^{\prime}}\left\|\mathbf{x}_{r}-\mathbf{{x}_{0}}^{\prime}\right\|_{F}^{2},$ (11)
@@
4 Experiments
@@
ISTD [41] Input Image 22.40 0.936 32.10 27.32 0.976 7.09 20.56 0.893 10.88
Guo et al. [16] 27.76 0.964 18.65 26.44 0.975 7.76 23.08 0.919 9.26
ST-CGAN [41] 33.74 0.981 9.99 29.51 0.958 6.05 27.44 0.929 6.65
DHAN [4] 35.53 0.988 7.49 31.05 0.971 5.30 29.11 0.954 5.66
Fu et al. [9] 34.71 0.975 7.91 28.61 0.880 5.51 27.19 0.945 5.88
DC-ShadowNet [20] 31.69 0.976 11.43 28.99 0.958 5.81 26.38 0.922 6.57
Zhu et al. [50] 36.95 0.987 8.29 31.54 0.978 4.55 29.85 0.960 5.09
ShadowFormer [14] 38.19 0.991 5.96 34.32 0.981 3.72 32.21 0.968 4.09
ShadowDiffusion [15] 40.15 0.994 4.13 33.70 0.977 4.14 32.33 0.969 4.12
@@
SRD [31] Input Image 18.96 0.871 36.69 31.47 0.975 4.83 18.19 0.830 14.05
Guo et al. [16] - - 29.89 - - 6.47 - - 12.60
DeshadowNet [31] - - 11.78 - - 4.84 - - 6.64
DHAN [4] 33.67 0.978 8.94 34.79 0.979 4.80 30.51 0.949 5.67
Fu et al. [9] 32.26 0.966 9.55 31.87 0.945 5.74 28.40 0.893 6.50
DC-ShadowNet [20] 34.00 0.975 7.70 35.53 0.981 3.65 31.53 0.955 4.65
BMNet [49] 35.05 0.981 6.61 36.02 0.982 3.61 31.69 0.956 4.46
SG-ShadowNet [40] - - 7.53 - - 2.97 - - 4.23
Zhu et al. [50] 34.94 0.980 7.44 35.85 0.982 3.74 31.72 0.952 4.79
ShadowFormer [14] 36.91 0.989 5.90 36.22 0.989 3.44 32.90 0.958 4.04
ShadowDiffusion [15] 38.47 0.987 4.98 37.78 0.985 3.44 34.73 0.957 3.63
@@
ISTD+ [22] Input Image 18.96 0.871 36.69 31.47 0.975 4.83 18.19 0.830 14.05
ST-CGAN [41] - - 13.40 - - 7.70 - - 8.70
DeshadowNet [31] - - 15.90 - - 6.00 - - 7.60
Fu et al. [9] 36.04 0.976 6.60 31.16 0.876 3.83 29.45 0.840 4.20
DC-ShadowNet [20] 32.20 0.976 10.43 33.21 0.963 3.67 28.76 0.922 4.78
BMNet [49] 37.30 0.990 6.19 35.06 0.974 3.09 32.30 0.955 3.60
SG-ShadowNet [40] - - 5.9 - - 2.90 - - 3.4
ShadowFormer [14] 39.67 - 5.20 38.82 - 2.30 35.46 - 2.80
ShadowDiffusion [15] 39.82 - 4.90 38.90 - 2.30 35.68 0.970 2.70
@@
Table 1: The quantitative results of shadow removal using our model and recent methods on ISTD [41], SRD [31] and ISTD+ [22] datasets. ‘-’ represents the values that were not reported in the original paper.
4.1 Experimental Setups
Datasets. We used three standard benchmark datasets for our shadow removal experiments: 1) ISTD dataset [41] includes 1,330 training and 540 testing triplets (shadow images, masks and shadow-free images); 2) Adjusted ISTD (ISTD+) dataset [22] reduces the illumination inconsistency between the shadow and shadow-free image of ISTD using linear regression; 3) SRD dataset [31] consists of 2,680 training and 408 testing pairs of shadow and shadow-free images without the ground-truth shadow masks. Referring to ShadowDiffusion [15], we applied the prediction mask provided by DHAN [4] for the training and testing.
Figure 4: Qualitative comparison with DC-ShadowNet [20], Fu et al. [9], BMNet [49], ShadowFormer [14] and ShadowDiffusion [15] on ISTD+ [22]. The values represent the ‘PSNR/SSIM’. The best and second-best results are illustrated in red and blue.
Implementation Details. Our Diff-Shadow is implemented by PyTorch, which is trained using eight NVIDIA GTX 2080Ti GPUs. The Adam optimizer [21] is applied to optimize the parallel UNets with a fixed learning rate of $lr=2e^{-4}$ without weight decay and the training epoch is set as 2,000. The exponential moving average [28] with a weight of 0.999 is applied for parameter updating.
Training Specifications. For each dataset, the images are resized into $256\times 256$ for training. For each iteration, we initially sampled 8 images from the training set and randomly cropped 16 patches with size $64\times 64$ from each image plus a down-sampled global image with the same size corresponding to each patch. This process resulted in mini-batches consisting of 96 samples. We used input time step embeddings for $t$ through sinusoidal positional encoding [39] and provided these embeddings as inputs to each residual block in the local and global branches, allowing the model to share parameters across time. During sampling, the patch size $R$ is set to 64 and the step size $r$ is set to 8 for covering the whole image.
Evaluation Measures. Following previous works [16, 31, 41, 4, 9, 20, 49, 14, 15], we utilized the Root Mean Square Error (RMSE) in the LAB color space as the quantitative evaluation metric of the shadow removal results, comparing to the ground truth shadow-free images. Lower RMSE scores correspond to better performance. In addition, we also adopt the common Peak Signal-to-Noise Ratio (PSNR) and the Structural Similarity Index Measure (SSIM) [42] to measure the performance in the RGB color space.
4.2 Comparison with State-of-the-art Methods
@@
We compare the proposed Diff-Shadow with the popular or state-of-the-art (SOTA) shadow removal algorithms, including one traditional method: Guo et al. [16], several deep learning-based methods: DeshadowNet [31], ST-CGAN [41], DHAN [4], Fu et al. [9], DC-ShadowNet [20], BMNet [49], ShadowFormer [14], and one diffusion-based method: ShadowDiffusion [15]. The shadow removal results of the competing methods are quoted from the original papers or reproduced using their official implementations. Following the previous methods [14, 15], we evaluate the shadow removal results with $256\times 256$ resolution.
@@
Quantitative Evaluation. Table 1 shows the quantitative results on the test sets over ISTD [41], SRD [31], and ISTD+ [22], respectively. The superiority of the proposed method over competing approaches is evident across all three datasets, as it consistently achieves superior performance. Thanks to the powerful modeling capabilities of the diffusion model for images, the proposed method improves the PSNR of the SRD [31] dataset from 32.90dB (ShadowFormer) to 34.93dB (Ours) when compared with the most recent transformer-based method ShadowFormer. Furthermore, compared to ShadowDiffusion which is also a diffusion-based method, our parallel network architecture interacts the local and global information more effectively. The novel Global-guided Sampling Strategy (GSS) further enhances the accuracy of shadow-free image restoration, which results in an improvement in PSNR from 32.33dB (ShadowDiffusion) to 33.69dB (Ours) on the ISTD [41] dataset. In summary, our Diff-Shadow provides better shadow removal performance than any of the recently proposed models, achieving new state-of-the-art results on three widely-used benchmarks.
Methods S NS ALL Sampling S NS ALL
(a) Single-head(w/o global) 2.64 3.41 3.36 (e) w/o GSS 2.97 3.63 3.81
(b) Single-head(w global) 2.55 3.19 3.04 (f) $W^{d}=1$ 2.64 3.42 3.27
(c) Ours (w/o RCA) 2.31 2.14 2.73 (g) $S_{\text{bri}}$ 2.16 2.99 2.96
(d) Ours (complete model) 1.81 1.65 2.52 (h) $S_{\text{bri}}+S_{\text{m}}$ 1.81 1.65 2.52
Table 2: Ablation studies on SRD dataset [31]. We report RMSE values.
(a) (b) (c) (d) (e) (f)
RCA first-layer w/o second-layer first&mid-layer w/o first-layer
CA w/o mid-layer mid-layer w/o every-layer mid-layer
[…]
Table 3: Ablation studies to evaluate the effectiveness of different network structures on the ISTD+ [22] dataset.
@@
Qualitative Evaluation. We present the qualitative comparison in Fig. 4 and the supplementary file. The proposed Diff-Shadow stands out among previous methods for its remarkable ability to remove shadows, while preserving the surrounding image textures and luminance, even in areas of high contrast that pose challenges for shadow removal. For example, in the third example of Fig. 4, there is a distinct color disparity between the ground and the shadow. Previous methods struggled to successfully remove the fragmented shadows of tree branches at this location, resulting in uneven color distribution or incomplete removal results. The result of our method is the most similar to the ground truth and surpasses the second-best approach by nearly 7dB. This further demonstrates the superiority of the proposed method in perceiving the shadow and non-shadow illumination consistency.
4.3 Ablation Studies
Effectiveness of Parallel UNets. The global branch is first removed to only accommodate the local inputs. Consequently, the network degrades to a single-head UNet structure. As in Table 2 (a), the single-head UNet shows the worst performance due to missing global context information. We tried to concatenate the patch and global images together at the network input level, and the results in Table 2 (b) show that the introduction of global context information works better than patch input alone. Then, in order to verify the effectiveness of the RCA module, we retained the two-branch network mechanism but eliminated the RCA module from the first layer and kept only the concatenate connection in the middle layer. As can be seen from Table 2 (c), the RCA module is effective in interacting two branches and enhancing the final result.
@@ 4.3 Ablation Studies
Figure 6: Qualitative comparison with original resolution for BMNet [49], ShadowFormer [14], and ours on ISTD [22] dataset.
@@
Effectiveness of Global-guided Sampling Strategy. Table 2 (e) presents the experiments without overlapping between patches. It turns out significant boundaries between patches in Fig. 5 (a). Second, we established the adaptive noise weights as fixed $W^{d}=1$ in Table 2 (f), which leads to a degradation of the sampling procedure. Reflected in Fig. 5 (b), we can see that there is a large discrepancy between the results and the ground truth. Third, we controlled the noise merge weights solely by the difference between the recovered patch brightness and the recovered global image brightness, i.e., only brightness-based score $S_{\text{bri}}$ exists. The qualitative patch image restored at the beginning of the sampling phase is heavily noisy in Fig. 5 (c). Therefore, it is not accurate to control the weights using brightness alone. The results in Table 2 (g) are inferior to those of our full strategy in Table 2 (h), which considers both the proportion of shaded areas in the patch and the brightness difference between the restored global and local image.
[…]
original size $640\times 480$ Input Image 20.33 0.874 11.35
DHAN [4] 27.88 0.921 6.29
BMNet [49] 29.02 0.928 4.17
ShadowFormer [14] 30.47 0.935 4.79
Ours 30.58 0.938 3.69
Table 4: Quantitative results on images with original resolution from ISTD [41].
Effectiveness of RCA module. We assess the impact of the RCA module and the Convolutional Aggregation (CA) at different layers on the ISTD+ [22] dataset.
Table 3 (a) and (b) demonstrate the lower performance when only the RCA module or only the CA module is applied. Table 3 (c) highlights the importance of the RCA module, showing its effectiveness in shallow feature fusion because low-level features capture essential elements, such as edges, colors, and corners, which are crucial for shadow removal tasks. Table 3 (d) confirms that adding the RCA module to both shallow and intermediate layers leads to unsatisfactory results due to limited mask size. Additionally, intermediate layer features primarily capture global semantic information, making them more suitable for fusion using the CA module. Table 3 (e) verifies the overall inferiority of solely incorporating the CA module at each layer compared to the full model in Table 3 (f) with the added RCA module. Therefore, the optimal approach is to apply the RCA module in the initial layer and the CA module in the intermediate layer.
4.4 Extension to Other Image Resolutions
Thanks to the patch-based diffusion modeling approach and the sampling process across overlapping patches during inference, the proposed method enables size-agnostic image restoration. As shown in Fig. 6, our Diff-Shadow maintains excellent illumination consistency even at the original image resolution ($640\times 480$) of the ISTD [22] dataset. Furthermore, as illustrated in Table 4, the proposed Diff-Shadow also achieves state-of-the-art performances. Note that the comparison methods are trained on images of original resolution, while our training is based on images with a resolution of $256\times 256$.
5 Conclusion

### v2

@@ Introduction
$\frac{1}{\sum_{i}W^{d_{i}}}\sum_{i=1}^{N_{p}}\left(W^{d_{i}}\!\cdot\!\epsilon_{\theta}\left(\mathbf{x}_{t}^{(i)},{\mathbf{x}_{L}}^{(i)},{\mathbf{x}_{G}},t\right)\right),$ (8)
@@
Optimization
The Diff-Shadow involves two parallel UNet networks operating the local patch optimization and the global shadow-free image reconstruction. The objective function includes the following two items: the diffusive objective function $L_{\text{diff}}$ and the global loss function $L_{\text{global}}$. Total loss $\mathcal{L}_{\text{total }}=\mathcal{L}_{\text{diff }}+\lambda\mathcal{L}_{\text{global}}$, where $\lambda$ is a hyper-parameter to balance the contributions of the two losses. We empirically set it to 1.
The Diffusive Objective Function. Refer to (Ho, Jain, and Abbeel 2020), the $\mathcal{L}_{\text{diff}}$ is defined as:
$\mathcal{L}_{\text{diff}}=\mathbb{E}_{\mathbf{x}_{0},t,\boldsymbol{\epsilon}}\left\|\boldsymbol{{\epsilon}}_{t}-\boldsymbol{\epsilon}\right\|_{F}^{2}.$ (9)
@@
$\mathcal{L}_{\text{global}}=\mathbb{E}_{\mathbf{{x}_{0}}^{\prime},\mathbf{{x}_{m}}^{\prime}}\left\|\mathbf{x}_{r}-\mathbf{{x}_{0}}^{\prime}\right\|_{F}^{2},$ (10)
@@
Experiments
Experimental Setups
Datasets. We used three standard datasets: 1) ISTD dataset (Wang, Li, and Yang 2018) includes 1,330 training and 540 testing triplets (shadow images, masks and shadow-free images); 2) Adjusted ISTD (ISTD+) dataset (Le and Samaras 2019) reduces the illumination inconsistency between the shadow and shadow-free image of ISTD; 3) SRD dataset (Qu et al. 2017) consists of 2,680 training and 408 testing pairs of shadow and shadow-free images without the shadow masks. Referring to ShadowDiffusion, we applied the prediction mask provided by DHAN (Cun, Pun, and Shi 2020) for the training and testing.
Implementation Details. Our Diff-Shadow is trained using eight NVIDIA GTX 2080Ti GPUs. The Adam optimizer (Kingma and Ba 2015) is applied to optimize the parallel UNets with a fixed learning rate of $lr=2e^{-4}$ without weight decay and the training epoch is set as 2,000. The exponential moving average (Nichol and Dhariwal 2021) with a weight of 0.999 is applied for parameter updating. Further details regarding the inference time and network complexity can be found in the supplementary.
Training Specifications. For each dataset, the images are resized into 256 $\times$ 256 for training. For each iteration, we initially sampled 8 images from the training set and randomly cropped 16 patches with size 64 $\times$ 64 from each image plus a down-sampled global image with the same size corresponding to each patch. This process resulted in mini-batches consisting of 96 samples. We used input time step embeddings for $t$ through sinusoidal positional encoding (Vaswani et al. 2017) and provided these embeddings as inputs to each residual block in the local and global branches, allowing the model to share parameters across time. During sampling, the patch size $R$ is set to 64 and the step size $r$ is set to 8 for covering the whole image.
Evaluation Measures. Following previous works (Fu et al. 2021; Jin, Sharma, and Tan 2021; Zhu et al. 2022a; Guo et al. 2023a, b), we utilized the Root Mean Square Error (RMSE) in the LAB color space as the quantitative evaluation metric of the shadow removal results, comparing to the ground truth shadow-free images. Lower RMSE scores correspond to better performance. In addition, we also adopt the common Peak Signal-to-Noise Ratio (PSNR) and the Structural Similarity Index Measure (SSIM) (Wang et al. 2004) to measure the performance in the RGB color space.
@@
ISTD Input Image 22.40 0.936 32.10 27.32 0.976 7.09 20.56 0.893 10.88
Guo et al. 27.76 0.964 18.65 26.44 0.975 7.76 23.08 0.919 9.26
ST-CGAN 33.74 0.981 9.99 29.51 0.958 6.05 27.44 0.929 6.65
DHAN 35.53 0.988 7.49 31.05 0.971 5.30 29.11 0.954 5.66
Fu et al. 34.71 0.975 7.91 28.61 0.880 5.51 27.19 0.945 5.88
DC-ShadowNet 31.69 0.976 11.43 28.99 0.958 5.81 26.38 0.922 6.57
Zhu et al. 36.95 0.987 8.29 31.54 0.978 4.55 29.85 0.960 5.09
ShadowFormer 38.19 0.991 5.96 34.32 0.981 3.72 32.21 0.968 4.09
ShadowDiffusion 40.15 0.994 4.13 33.70 0.977 4.14 32.33 0.969 4.12
@@
SRD Input Image 18.96 0.871 36.69 31.47 0.975 4.83 18.19 0.830 14.05
Guo et al. - - 29.89 - - 6.47 - - 12.60
DeshadowNet - - 11.78 - - 4.84 - - 6.64
DHAN 33.67 0.978 8.94 34.79 0.979 4.80 30.51 0.949 5.67
Fu et al. 32.26 0.966 9.55 31.87 0.945 5.74 28.40 0.893 6.50
DC-ShadowNet 34.00 0.975 7.70 35.53 0.981 3.65 31.53 0.955 4.65
BMNet 35.05 0.981 6.61 36.02 0.982 3.61 31.69 0.956 4.46
SG-ShadowNet - - 7.53 - - 2.97 - - 4.23
Zhu et al. 34.94 0.980 7.44 35.85 0.982 3.74 31.72 0.952 4.79
ShadowFormer 36.91 0.989 5.90 36.22 0.989 3.44 32.90 0.958 4.04
ShadowDiffusion 38.47 0.987 4.98 37.78 0.985 3.44 34.73 0.957 3.63
@@
ISTD+ Input Image 18.96 0.871 36.69 31.47 0.975 4.83 18.19 0.830 14.05
ST-CGAN - - 13.40 - - 7.70 - - 8.70
DeshadowNet - - 15.90 - - 6.00 - - 7.60
Fu et al. 36.04 0.976 6.60 31.16 0.876 3.83 29.45 0.840 4.20
DC-ShadowNet 32.20 0.976 10.43 33.21 0.963 3.67 28.76 0.922 4.78
BMNet 37.30 0.990 6.19 35.06 0.974 3.09 32.30 0.955 3.60
SG-ShadowNet - - 5.9 - - 2.90 - - 3.4
ShadowFormer 39.67 - 5.20 38.82 - 2.30 35.46 - 2.80
ShadowDiffusion 39.82 - 4.90 38.90 - 2.30 35.68 0.970 2.70
@@
Table 1: The quantitative comparisons on ISTD, SRD, and ISTD+ datasets. ‘-’ represents the values that were not reported.
Figure 4: Qualitative comparisons on ISTD+. The values represent the ‘PSNR/SSIM’.
Comparison with the State-of-the-art
@@
We compare the proposed Diff-Shadow with the popular or state-of-the-art shadow removal algorithms, including one traditional method: Guo et al. (Guo, Dai, and Hoiem 2012), several deep learning-based methods: DeshadowNet (Qu et al. 2017), ST-CGAN (Wang, Li, and Yang 2018), DHAN (Cun, Pun, and Shi 2020), Fu et al. (Fu et al. 2021), DC-ShadowNet (Jin, Sharma, and Tan 2021), BMNet (Zhu et al. 2022a), ShadowFormer (Guo et al. 2023a), and one diffusion-based method: ShadowDiffusion (Guo et al. 2023b). The shadow removal results of the competing methods are quoted from the original papers. Following the previous methods (Guo et al. 2023a, b), we evaluate the removal results with 256 $\times$ 256 resolution, while the results for other resolutions are detailed in the supplementary.
@@
Quantitative Evaluation. Table 1 shows the quantitative results on the test sets over ISTD, SRD, and ISTD+, respectively. The superiority of the proposed method over competing approaches is evident across all three datasets, as it consistently achieves superior performance. Thanks to the powerful modeling capabilities of the diffusion model for images, the proposed method improves the PSNR of the SRD from 32.90 dB to 34.93 dB when compared with the transformer-based method ShadowFormer. Furthermore, compared to ShadowDiffusion which is also a diffusion-based method, our parallel network architecture interacts the local and global information more effectively. The novel Global-guided Sampling Strategy (GSS) further enhances the accuracy of shadow-free image restoration, which results in an improvement in PSNR from 32.33 dB to 33.69 dB on the ISTD. In summary, our Diff-Shadow achieves new state-of-the-art results on three widely-used benchmarks.
@@
Qualitative Evaluation. We present the qualitative comparison in Fig. 4 and the supplementary. Diff-Shadow stands out among previous methods for its ability to remove shadows, while preserving the surrounding image textures and luminance, even in areas of high contrast that pose challenges for shadow removal,  e.g., in the third example of Fig. 4, there is a distinct color disparity between the ground and the shadow. Previous methods struggled to remove the fragmented shadows of branches at this location, resulting in uneven color distribution or incomplete removal results. The result of our method is the most similar to the ground truth.
Methods S NS ALL
(a) Single-head(w/o global) 2.64 3.41 3.36
(b) Single-head(w global) 2.55 3.19 3.04
(c) Ours (w/o RCA) 2.31 2.14 2.73
(d) Ours (complete model) 1.81 1.65 2.52
Table 2: Ablation studies of different network structures on SRD dataset. We report RMSE values.
Sampling S NS ALL
(a) w/o GSS 2.97 3.63 3.81
(b) $W^{d}=1$ 2.64 3.42 3.27
(c) $S_{\text{bri}}$ 2.16 2.99 2.96
(d) $S_{\text{bri}}+S_{\text{m}}$ 1.81 1.65 2.52
Table 3: Ablation studies of different sampling strategies on SRD dataset. We report RMSE values.
@@
Ablation Studies
Effectiveness of Parallel UNets. The global branch is removed to only accommodate the local inputs. Consequently, the network degrades to a single-head UNet structure. As in Table 2 (a), the single-head UNet shows the worst performance due to missing global context information. We tried to concatenate the patch and global images together at the network input level, and the results in Table 2 (b) show that the introduction of global context information works better than patch input alone. Then, to verify the effectiveness of the RCA module, we retained the two-branch network mechanism but eliminated the RCA module from the first layer and kept only the concatenate connection in the middle layer. As shown in Table 2 (c), the RCA module is effective in interacting two branches and enhancing the final result.
@@
Effectiveness of Global-guided Sampling Strategy. Table 3 (a) presents the experiments without overlapping between patches. It turns out significant boundaries between patches in Fig. 5 (a). Second, we established the adaptive noise weights as fixed $W^{d}=1$ in Table 3 (b), which leads to a degradation of the sampling procedure. Reflected in Fig. 5 (b), we can see that there is a large discrepancy between the results and the ground truth. Third, we controlled the noise merge weights solely by the difference between the recovered patch brightness and the recovered global image brightness, i.e., only brightness-based score $S_{\text{bri}}$ exists. The qualitative patch image restored at the beginning of the sampling phase is heavily noisy in Fig. 5 (c). Therefore, it is not accurate to control the weights using brightness alone. The results in Table 3 (c) are inferior to those of our full strategy in Table 3 (d), which considers both the proportion of shaded areas in the patch and the brightness difference between the restored global and local image.
Conclusion